Journal of Lie Theory Volume 11 (2001) 255 \begin{matrix}c \\ ◯ \end{matrix} 2001 Heldermann Verlag Berlin

Erratum to “Berezin transform on line bundles over bounded symmetric domains”

Genkai Zhang

Communicated by K.-H.Neeb

The text of Remark 4.10 on p.123 in my article entitled “Berezin transform on line bundles over bounded symmetric domains,” Volume 10 (2000), 111–126, is incorrect as it stands. It should read: Theorem 4.6 is still valid under weaker conditions on ν and σ such as indicated in Remark 3.2 above.

Genkai Zhang Department of Mathematics Chalmers University of Technology and Göteborg University S-412 96 Göteborg Sweden genkai@math.chalmers.se

Received May 29, 2000

ISSN 0949–5932 / $2.50 \begin{matrix}c \\ ◯ \end{matrix} Heldermann Verlag